Find the electric flux in a sphere of radius 0.50 cm from a charge of 60 nC


1
Expert's answer
2022-03-18T09:08:21-0400

The electric flux is given by

"\Phi=\oint EdA=\frac{q}{\epsilon_0}"

"\Phi=\frac{60\times 10^{-9}}{8.85\times 10^{-12}}=6.8\times 10^3\:\rm N\cdot m^2/C"